How many solutions does the equation |x+7|=1 have?
Answer by jim_thompson5910(35256)

|x+7|=1

x+7=1 or x+7=-1

x=1-7 or x=-1-7

x = -6 or x = -8

So there are 2 solutions (and they are x = -6 or x = -8)

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------